# Fluid flow inside slit-shaped nanopores: the role of surface morphology at the molecular scale

### v1

@@ II.1 Hagen–Poiseuille flow in a slit-shaped conduit
At low Reynolds number, the l.h.s. of Eq. 1 is negligible and the velocity field is aligned with the direction of the mean flow, ${\boldsymbol{v}}=(v_{x},0,0)$. Due to the incompressibility of the fluid, low Mach number, and in the absence of an external force, the pressure gradient is constant: $\partial_{x}P(x,z)=-\Delta P/L$. Thus, the problem becomes $x$-independent and the $x$-component of the flow field obeys:
$0=\Delta P/L+\eta\,\partial^{2}_{z}v_{x}(z)\,.$ (3)
Applying the boundary condition in Eq. 2 results in the Hagen–Poiseuille flow profile [79, 80]:
$v_{x}(z)=\frac{\Delta P}{2\eta L}\left(\frac{1}{4}D^{2}-z^{2}\right).$ (4)
@@
Unlike in macroscale systems, where the friction near surfaces is fully developed, at the atomistic scale, the fluid molecules slip along the surface and, the arrangement of surface atoms, as well as their interaction with the fluid molecules, affects the boundary layer of the flow. These lead to an effective full or partial slip of the flow field, described by Navier’s partial slip condition [77, 81]:
@@
where $\beta_{\text{slip}}$ is the slip coefficient, representing the distance below the surface at which the no-slip condition would hold if the velocity was linearly extrapolated into the wall. As an alternative to $\beta_{\text{slip}}$, we introduce the slip length ${\delta_{\text{slip}}}$ as an apparent displacement of the pore surface from, e.g., $z=D/2$ to $z=D/2+{\delta_{\text{slip}}}$. The velocity profile for partial slip follows from solving Eq. 3 for the boundary condition (5) instead of the no-slip condition Eq. 2:
$v_{x}(z)=\frac{\Delta P}{2\eta L}\left(\frac{1}{4}(D+2{\delta_{\text{slip}}})^{2}-z^{2}\right);$ (6)
in this case, the two definitions of the slip length are related via:
@@
$J_{\varrho}=\frac{1}{D}\int_{-D/2}^{D/2}\varrho v_{x}(z)\,\mathrm{d}z\,,$ (8)
which is given as the momentum density field, $\varrho{\boldsymbol{v}}$, averaged over the cross section of the conduit. In the case of Hagen–Poiseuille flow without slip, one finds:
$J_{\varrho}=\frac{2}{3}\varrho v_{x}(0)=\frac{D^{2}\Delta P}{12\nu L}\,,$ (9)
@@
using Eq. 4 and the kinematic viscosity $\nu=\eta/\varrho$. In the field of porous media, the proportionality between fluid flow and the applied pressure gradient is known as Darcy’s law[82, 48], which defines the permeability $k$ of the medium:
$J_{\varrho}=\frac{k\Delta P}{\nu L}\,.$ (10)
Comparison with Eq. 9 yields the permeability of the slit-shaped conduit:
$k=\frac{D^{2}}{12}.$ (11)
If slip at the surface is taken into account, one calculates from Eq. 6:
$k$ $=\frac{D^{2}}{12}+\beta_{\text{slip}}(D/2)$ (12)
$=\frac{1}{12}\left(D^{2}+6D{\delta_{\text{slip}}}+6{\delta_{\text{slip}}}^{2}\right),$ (13)

### v2

@@ III.1 Poiseuille flow in a slit-shaped conduit
The steady flow of a uniform, incompressible, Newtonian fluid at low Reynolds number is governed by the stationary Stokes equation for the velocity field ${\boldsymbol{v}}(x,z)$ and the pressure field $P(x,z)$. In the absence of an external force and far from the pore entrance, the problem is $x$-independent and the flow is fully developed and aligned with the axis of the conduit, ${\boldsymbol{v}}=(v_{x},0,0)$. For low Mach number, due to the incompressibility, the pressure gradient is aligned with the flow and thus constant, $\partial_{x}P(x,z)=-\Delta P/L$. This simplifies the stationary Stokes equation to read:
$0=\Delta P/L+\eta\partial^{2}_{z}v_{x}(z),$ (3)
where $\eta$ is the dynamic shear viscosity. In Poiseuille theory, $\eta$ is assumed to be a constant parameter.
@@
At macroscopic scales, no-slip boundary conditions hold at the surfaces of the conduit, ${\boldsymbol{v}}(x,\pm D/2)=0$. At the atomistic scale, the friction near surfaces is not fully developed and fluid molecules can slip along the surface. The arrangement of surface atoms and also their interaction with the fluid molecules affect the boundary layer of the flow. These lead to an effective full or partial slip of the flow field, described by Navier’s partial slip condition [86, 85]:
@@
where ${\beta_{\text{slip}}}$ is the slip length, representing the distance below the surface at which the no-slip condition would hold if the velocity was linearly extrapolated into the wall. As an alternative, one may use a parabolic extrapolation of the flow profile near the surface, which yields a length ${\delta_{\text{slip}}}$ that is interpreted as an apparent displacement of the pore surface from, e.g., $z=D/2$ to $z=D/2+{\delta_{\text{slip}}}$. Both lengths are related via:
@@
Solving Eq. 3 for the boundary condition (4) leads to the velocity profile of Poiseuille flow:
$v_{x}(z)=\frac{\Delta P}{2\eta L}\left[\left(\frac{D}{2}+{\delta_{\text{slip}}}\right)^{2}-z^{2}\right].$ (6)
The no-slip solution is contained for ${\delta_{\text{slip}}}={\beta_{\text{slip}}}=0$.
III.2 Permeability
@@
$J_{\varrho}=\frac{1}{D}\int_{-D/2}^{D/2}m\varrho v_{x}(z)\,\mathrm{d}z\,,$ (7)
@@
given as the momentum density field, $m\varrho{\boldsymbol{v}}$, averaged over the cross section of the conduit. The fluid flow is proportional to the applied pressure gradient, if the latter is not too large. This is known as Darcy’s law in the field of porous media [87, 48], which defines the permeability $k$ of the medium:
$J_{\varrho}=\frac{k\Delta P}{\nu L}$ (8)
in terms of the kinematic viscosity $\nu=\eta/(m\varrho)$. For Poiseuille flow, Eq. 6, one finds for the permeability of the slit-shaped conduit:
$k$ $=\frac{D^{2}}{12}+{\beta_{\text{slip}}}\,\frac{D}{2}\,.$ (9)